OpenStudy (anonymous):

how do you find the k value for S= 500,000/1+0.6e^kt

OpenStudy (mathteacher1729):

You need to know what S is otherwise you can't solve for k. Additionally, you'd need to know what t is as well. You can solve one equation in one variable. One equation in THREE variables is not solvable. :(

OpenStudy (anonymous):

S(1+.6e^kt)=500000

OpenStudy (anonymous):

I think that thing goes to zero as t approaches infinity

OpenStudy (anonymous):

hey i feel u r making a mistake

OpenStudy (anonymous):

do u feel it is correct in financial sense???????

OpenStudy (anonymous):

when you have multiple variables, it's all about perspective. You can fix different variables and manipulate others to see what it would be like

OpenStudy (anonymous):

indy it was not for u u must have calculated the right value my question is for dria

OpenStudy (anonymous):

@ Dria do u feel it is correct question

OpenStudy (mathteacher1729):

Here are the steps: Go to http://www.wolframalpha.com and type solve 500,000/(1+0.6e^kt) = S for k Click "show steps".

OpenStudy (anonymous):

\[YTM = \frac{{\frac{FV-P}{n}} + A}{\frac{FV+P}{2}}\]

OpenStudy (anonymous):

ln (e^kt) = kt

OpenStudy (anonymous):

write the full question then only i can help u out!!

OpenStudy (anonymous):

@ indy here we are talking about Time Value of Money And in this case \[\frac{1}{(1+k)^t} = Exp^{-kt}\] \[Exp^{kt} = (1+K)^t\]

OpenStudy (anonymous):

hey dria u dre

OpenStudy (anonymous):

what's your point?

OpenStudy (anonymous):

u can see Dria has asked 500000/(1 +0.6e^kt) which is wrong in financial terms so i m trying to tell her that exponential sign will not be there

OpenStudy (anonymous):

okay, but your formula is not necessary. You can solve this problem outright with the necessary parameters.

OpenStudy (anonymous):

yeah u r ryt but after assuming a number of things right but here u have to calculate the yield rate of a bond in that case how u will??????

OpenStudy (anonymous):

im a little confused

OpenStudy (anonymous):

u just tell me the whole problem then will help u@ DRIA

OpenStudy (anonymous):

Dria, this problem is missing information

OpenStudy (anonymous):

yes Indy now u got my point

OpenStudy (anonymous):

the question said in the year of 1994 t=0 and in 1996 S=300,000 which would make t=2 but it asked for the value of k how do i solve for k that is the only missing info

OpenStudy (anonymous):

I am very familiar with the concepts. There are certain structural properties of the equation that she presented, of which, I was trying to describe.

OpenStudy (anonymous):

yeah i know actually it is in finance

OpenStudy (anonymous):

that's not the only place that equations like this show up.

OpenStudy (anonymous):

one thing more from where did u got 0.6 in ur question?

OpenStudy (anonymous):

\[\s=500,000\div1+0.6e^{kt}\]

OpenStudy (anonymous):

no i dint meant it only comes in Finance........but this type of problems are their in finance

OpenStudy (anonymous):

that was apart of the question in the worksheet

OpenStudy (anonymous):

\[300000 = \frac{500000}{1+0.6e^{kt}}\]

OpenStudy (anonymous):

you will still need t

OpenStudy (anonymous):

i said for the year that s=300,000 t was equal to 2

OpenStudy (anonymous):

so then, 300000+300000(.6)e^2k = 500000

OpenStudy (anonymous):

can you use a calculator?

OpenStudy (anonymous):

k comes out to be 5.2%

OpenStudy (anonymous):

e^2k = 1.111111 now solve for k

OpenStudy (anonymous):

what is the answer to this problem

OpenStudy (anonymous):

DRIA?????????

OpenStudy (anonymous):

im trying to solve it with my calculator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the answer is correct thank you

OpenStudy (anonymous):

ok sorry i am trying to apply the answer to the equation thanks sheg for being patient with me

OpenStudy (anonymous):

k i will again thanks

OpenStudy (anonymous):

i dnt have fb lol

OpenStudy (anonymous):

ok gud bye

OpenStudy (anonymous):

bye bye